Fermi National Laboratory
 

Accelerator Update for March 5, 2007 - March 7, 2007

 

Monday March 5

The day shift began with Operations monitoring the Recycler stash (300.0E10), with the Antiproton Source (Pbar) not stacking (8.0E10) due to a target problem, with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations prepared the CDF and D0 collision halls for access at 8:20 AM.

At 8:33 AM, a Linac RF station (LRF2) tripped off. Experts investigated and reset the station.

Operations prepared Pbar for access at 9:35 AM to repair the target rotation problem.

Pbar experts reported at 10:17 AM that they found a LCW leak. LCW technicians were notified and they repaired the leak.

Pbar experts reported at 10:54 AM that they had completed their work on the target.

D0 came out of access at 10:57 AM.

Operations halted the NuMI beam briefly at 11:04 AM, so the experiment could prepare for their 11 batch slipstacking studies. The studies began about 30 minutes later.

Experts worked on a Tevatron (TeV) quadrupole power supply (T:QDD1) and discovered a loose connection on a transformer for the power supply’s phase circuitry. The experts reported at 11:19 AM that the loose connection was the cause of the phase loss trip on Saturday.

Operations established beam to the Meson MTest area for studies at 11:25 AM.

The Neutron Therapy Facility (NTF) reported at 12:09 PM that they treated their last patient for the day.

CDF came out of access at 12:14 PM.

The NuMI studies ended at 12:34 PM. NuMI resumed normal operations.

Operations put the TeV through a dry squeeze at 12:54 PM.

Operations put the TeV into shot setup at 1:30 PM.

At 2:57 PM, a Booster RF station (BRF6) tripped off. It reset.

At 3:19 PM, the NuMI beam permit tripped off due to a power supply’s current (QD108) drifting low. It reset.

Operations established store 5264 at 3:38 PM with an initial luminosity of 241.67E30.

Operations reported at 3:59 PM that the Recycler had stashed antiprotons once during the day shift.

Operations reported at 5:08 PM that MI experts were conducting barrier bucket studies and that Recycler experts were conducting Electron Coolingg studies.

At 5:29 PM, a Linac Klystron RF station (KRF1) tripped off on a LCW flow fault. While investigating, operators discovered that KRF5 had tripped off on reflective power. Both stations were brought back on line.

At 6:16 PM, Operations reported that the Pbar dry engine had tripped off. Operations contacted cryo system techs at CHL who called in an expert to repair the engine.

At 6:36 PM, NuMI tripped off due to a bad beam position on their target. It reset.

MI studies ended at 6:58 PM.

Cryo system techs reported at 7:57 PM that they had reset the Pbar dry engine drive controller and that the engine was running.

At 9:04 PM, KRF6 tripped off. It reset.

At 9:40 PM, LRF4 tripped off. An expert reset the station and ran it slowly up to its normal gradient.

At 11:39 PM, LRF4 tripped off. It reset.

At 11:40 PM, KRF5 tripped off. It reset.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Tuesday March 6

The midnight shift began with Operations monitoring store 5264 (89.83E30) the stash (119.5E10), with Pbar stacking (30.8E10), and with NuMI taking beam.

At 12:41 AM, BRF6 tripped off. It reset.

At 2:39 AM, BRF7 tripped off. It reset.

At 3:47 AM, BRF10 tripped off on a supply ground fault. It reset.

At 6:47 AM, there was a series of trips starting with LRF4 that including KRF1, 3, and 4 that lasted for about 30 minutes. Everything eventually reset.

Switchyard (SY) experts began taking study beam at 7:21 AM.

Operations reported at 7:56 AM that the Recycler had stashed three times during the midnight shift.

At 8:32 AM, LRF1 tripped off. It reset.

At 9:39 AM, MIRF14 tripped off. It reset.

At 10:03 AM, a TeV sector F17 dipole power supply (I:F17B3) tripped off. This problem halted beam to Pbar. (Beam transfers between Pbar and MI go through F sector.) Operations switched to a NuMI only timeline. An expert investigated and found a faulty pressure relief valve alarm; he bypassed the alarm. Beam to Pbar resumed at 10:24 AM.

At 12:26 PM, a Pbar Debuncher RF station (DRF1-7) tripped off and wouldn’t reset. An expert investigated, replaced a power supply, and had the station back up by 12:59 PM.

At 1:30 PM, the TeV leak alarm sounded. The sector A3 LCW supply pressure also alarmed. Operators and water techs searched the service buildings, but found no leaks.

Operations reported at 1:50 PM that by using humidity readings they found the LCW leak; it was in Pbar’s PreVault area. Operators halted stacking and then prepared the area for an access.

Operations reported at 2:35 PM that the Pbar LCW leak was on water line that fed D:H926. The LCW system lost about 970 gallons.

LCW techs reported at 3:16 PM that they had repaired the PreVault leak.

Pbar resumed stacking at 3:26 PM.

At 3:45 PM, LRF3 tripped off. An expert reset the station.

Operations reported at 4:07 PM that the Recycler had stashed twice during the day shift.

Meson experts completed their MTest beam studies at 5:12 PM.

Operations terminated store 5264 at 6:02 PM.

Operations put the TeV into shot setup at 6:16 PM.

At 7:26 PM, Operations reported problems with TeV proton coalescing. Operations contacted experts who reported that the problem was with cavity #3. They replaced a bad amplifier. Normal coalescing resumed at 9:05 PM.

At 9:14 PM, KRF3 tripped off. It reset.

Operations established store 5265 at 10:31 PM with an initial luminosity of 249.84E30.

At 11:57 PM, KRF5 tripped off. It reset.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Wednesday March 7

The midnight shift began with Operations monitoring store 5265 (181.06E30) and the stash (81.8E10), with Pbar stacking (32.7E10), and with NuMI taking beam.

At 12:13 AM, KRF3 tripped off. It reset.

At 1:57 AM, the MI beam permit tripped off for the first of five times in six hours due to the MI-8 collimators. The permit reset each time.

At 3:31 AM, NuMI tripped off due to bad beam positions. It reset.

At 4 AM, KRF5 tripped off. It reset.

At 4:49 AM, BRF7 & 8 tripped off due to a vacuum pump trip. Everything reset.

Operations reported at 7:55 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to repair a power supply LCW leak (V901) that just occurred and is holding off a stash. After the repair and stash, Pbar experts will conduct a four-hour reverse proton study. The next store will go in late this evening or early on the midnight shift. TeV experts will conduct an end-of-store study.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/07/2007    email Fermilab